# Canonicalization for Unreproducible Builds in Java

### v1

@@ 4.3.5. Unreproducible Versioning Properties
Git properties are embedded in Jar files in git.json or git.properties files. Maven plugin git-commit-id-maven-plugin  [^16] https://github.com/git-commit-id/git-commit-id-maven-plugin is commonly used to do that. These properties include timezone of commit, remote URL, number of git tags in repository, name of build host, time of build, name of builder, email of builder, branch name, number of commits, local branch name, and total commits in the repository. All of them sometimes diverge. For example, value of git.tags vary for Maven release org.apache.drill:drill-opentsdb-storage:1.21.0 due to additional number of tags in the repository during rebuild. Mitigation: The rebuilder should strip out the git properties from the Jar file before comparison or avoid comparison of the Git properties. Another solution is to fix the values for each git property. For example, total commits should be calculated up to the commit that is used to build the artifact. See issue on our repository  [^17] https://github.com/chains-project/reproducible-central/issues/19 to see more examples of differences in the git properties.
@@ 4.3.6. Unreproducible Timestamps
However, we observe that there are 10 other ways timestamps can be embedded in the JAR file. In our dataset, we observe that: 1) they can be embedded in the properties file section F.1, 2) generated documentation section F.2, 3) shell scripts section F.3, 4) executable binaries section F.4, 5) software bill of materials section F.5, 6) JVM bytecode section F.6, 7) file metadata section F.7, 8) MANIFEST.MF section F.8, and 9) NOTICE section F.9. 10) servlets created by Jasper JSP compiler section F.10. Mitigation: The solution proposed by Maven is partial, more engineering is needed to support project.build.outputTimestamp in all Maven plugins that create the above mentioned files.
@@ 4.3.7. Miscellaneous Unreproducibility Reasons
We report reasons of reproducibility that are too specific to be categorized in the above sections.
The first reason is due to the use of different build tools. For example, the Maven build tool is available as standalone tool and it also integrated into the Eclipse IDE. They embed information differently. m2e.projectLocation and m2e.projectName is embedded in pom.properties of org.spdx:spdx-maven-plugin:0.7.0 and org.owasp.antisamy:antisamy:1.7.3 section B.3. Even if the build tool is same, there are differences in rebuild due to usage of versions ranges of a plugin. For example, org.apache.aries:org.apache.aries.jax.rs.whiteboard:2.0.2 uses [1.2.5,) as the version range for flatten-maven-plugin in the POM file. This causes difference in the ouput of the plugin. Mitigation: The rebuilder should use the same build tool and its precise version as the builder. Moreover, the builder should avoid using version ranges for plugins in the POM file.
The next reason is due to a difference in the operating system. The reference version of org.apache.helix:helix-front:1.0.3 produces a Mach-O object file format (Mac) which is different from the ELF binary format produced by the rebuilt version (Linux). See https://github.com/chains-project/reproducible-central/issues/20#issuecomment-2583944777 for detailed diff. Mitigation: The builder should ensure that 1) the OS is specified in the build specification or 2) that the binary format is independent of the operating system used to build the artifact.
Usage of Universally Unique Identifier (UUID) is another reason of unreproducibility. We observe that Maven release org.apache.karaf.examples:karaf-docker-example-static-dist:4.4.0 uses UUID to generate the configuration files. Mitigation: The builder should use seed values to generate the UUID so that the UUID can be reproduced.
Maven releases org.apache.helix:helix-front:1.0.3, io.wcm:io.wcm.caconfig.editor:1.9.0, and org.apache.isis.viewer:isis-viewer-restfulobjects-viewer:2.0.0-M7 embed bundled JavaScript files in the Jar file. These files are subject to changes in variable names and new lines. Mitigation: The changes in variable names can be fixed by pinning the version of the bundler. For example, Goswami et al. (Goswami et al., 2020) show that UglifyJS can produce output with different variable names for the same input if the version of UglifyJS is different. The difference in new lines can be fixed by the rebuilder by canonicalizing them.
There is one instance where microbenchmarking results are embedded in the Jar file section G.4. Mitigation: The builder should not embed microbenchmarking results in the Jar file as they are solely for the purpose of improving performance (Costa et al., 2021) of the application and not needed for the end user.
Finally, there are some indices files created by tools like Jandex  [^19] https://smallrye.io/jandex/jandex/3.2.1/index.html - a space efficient Java class file indexer, EFXToolkit  [^20] https://github.com/OP-TED/efx-toolkit-java - a toolkit for expressing eForms business rules, and SpotlessFormatter  [^21] https://github.com/diffplug/spotless - a code formatter for Java. Mitigation: The builder should not embed these index files in the artifact as they are not required for the end user and they serve only as cache files for plugins that use them.
@@ 4.3.8. Mysteries
We observe changes in NOTICE, POM, and other configuration files that are not due to any of the above reasons and thus, we are unable to propose a mitigation strategy for them. Since these reasons of unreproducibility are so diverse, we discuss them one by one in appendix H.
{mdframed}
@@
We identify 6 main reasons of unreproducibility in Java and their mitigation: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. They represent the key challenges that researchers will address. For practitioners, they provide clear guidelines to diagnose and mitigate their unreproducible builds in Java.
@@ 5.2. Methodology
Chains-Rebuild  [^23] https://github.com/chains-project/chains-rebuild/commit/6dd67d5c7ac4db112f3419b5132d8f80a22cbe65 is our own fork of OSS-Rebuild, with crucial improvements dedicated to Maven reproducibility, which is unfortunately missing in the upstream project. We add support for canonicalizing build manifests and embedded versioning properties in JAR files as suggested in taxonomy (Table 1). For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file, a spurious change which is removed by Chains-Rebuild. For versioning properties, we remove the git.properties files embedded in the JAR file. Finally, we remove the pom.properties file.
We run OSS-Rebuild and Chains-Rebuild on the 12,283 reference and rebuild artifact pairs in order to analyze how effectively they make the build reproducible, that is when artifacts are bit-by-bit identical after canonicalization as shown below.
$canonicalize(ReferenceArtifact)=canonicalize(RebuildArtifact)$
@@ 5.3. Results
OSS-Rebuild (4ef4c01) 1165 (9.48%) 11118 (90.52%)
Chains-Rebuild (6dd67d5) 3036 (24.72%) 9247 (75.28%)
Table 2. Effectiveness of OSS-Rebuild and Chains-Rebuild on mitigating 12,283 unreproducible artifacts via canonicalization.
Table 2 shows the results of OSS-Rebuild and Chains-Rebuild on 12,283 unreproducible artifacts. The first column is the name of the tool and its commit hash we use to canonicalize the artifacts. The second column shows the number of artifact pairs that are successfully canonicalized by OSS-Rebuild and Chains-Rebuild. The third column shows the number of artifact pairs that are not properly canonicalized.
@@
Our publicly available prototype Chains-Rebuild can canonicalize (3036) 24.71% of the artifacts. Our study is the first large-scale experiment on artifact canonicalization, demonstrating that it is a promising mitigation of build unreproducibility in Maven.
@@ 6.2. Methodology
Bytecode canonicalization is a process of transforming the bytecode of a program into a representation that is independent of specific implementation details inserted by the compiler. This is either done by converting the bytecode to an intermediate form or by rewriting the bytecode. For example, subtraction of two integers is implemented as addition of a positive and negative integer in Java 6 and higher versions. Meanwhile, in Java 5 and lower versions, it is implemented as subtraction of two integers. Hence, if the program is compiled with Java 5 and 6, the two bytecode representations differ even though the program is semantically the same - performing subtraction of two integers. Canonicalization removes these differences by converting the bytecode to an intermediate form which is free from the implementation details that vary across different versions of the compiler.
@@
We run jNorm version 1.0.0 on the 898 reference rebuild artifact pairs that have differences in the bytecode. For maximum canonicalization, we run the tool with all the command line options as listed in the README 242424jNorm tool - https://github.com/stschott/jnorm-tool/tree/cec4645c5c9b52f73c347349bf14945b0eb55c87 of the repository and tool’s help output. As a result of canonicalizing the artifacts, jNorm produces files in the Jimple (Raja Vallée-Rai and Laurie J. Hendren, 1998) format. We save the canonicalized Jimple files for both the reference and the rebuild artifacts. There are cases where jNorm crashes and canonicalization fails, which we consider as a failure. We do not analyze them further as they are bugs in tool, but we upload the logs for future debugging. We run the Linux diff tool over the canonicalized pairs of Jimple files. If there is no difference, it means that canonicalization is a success, for reproducibility. In other words, successful canonicalization by jNorm produces two identical artifacts. We report the number of cases where the diff between the pair is empty. Finally, we analyze the cases where jNorm is unable to fully canonicalize the bytecode in order to understand its fundamental limitations and identify required future work.
@@ 6.3. Results
Out of 898 artifacts, jNorm’s execution leads to the following distribution: it successfully canonicalizes 267 (29.7%) artifacts which means it produces two identical artifacts; it leaves 478 (53.2%) artifacts with a diff after canonicalization; and there are 153 (17.1%) crashes, which are ignored.
@@
Next, we study the reasons why jNorm fails to produce identical artifacts when run on 478 cases. We categorize them based on the features in the bytecode that are not properly canonicalized by jNorm. Note that some unreproducible artifacts may fall into multiple categories as it is possible jNorm is unable to canonicalize multiple problems in the same artifact.
@@
    Listing 2: jNorm diff for control flow differnce.
@@
We note that multiple canonicalization tools can be used in conjunction. For instance, Chains-Rebuild and jNorm can be used together to canonicalize the JVM bytecode and the JAR files of a Maven artifact. Since Chains-Rebuild can canonicalize 3,036 artifacts and jNorm can canonicalize 267 artifacts, we can combine the 26.89% (3,303 / 12,283). Out of 3,303 artifacts, 2,578 artifacts are canonicalized such that 2,263 / 7,961 (28.43%) releases become fully reproducible.
{mdframed}
@@
jNorm is able to canonicalize 267 (29.7%) out of 898 bytecode artifacts. However, there are 478 (53.2%) artifacts that are still unreproducible after canonicalization which calls for improvements. We present 4 limitation of bytecode canonicalization related to structural change, control flow, embedded data, and optimization. Finally, our experiments demonstrate that combining artifact and bytecode canonicalization successfully mitigates unreproducibility in 26.89% (3,303 / 12,283) artifacts.
@@ 7.1. Reproducibility in Applications
The Reproducible Builds  [^30] https://reproducible-builds.org/ project works on the reproducibility of coreboot  [^31] https://tests.reproducible-builds.org/coreboot/coreboot.html - an open-source firmware for x86 and ARM systems, Debian  [^32] https://tests.reproducible-builds.org/debian/reproducible.html - a Linux distribution, FreeBSD  [^33] https://tests.reproducible-builds.org/freebsd/freebsd.html and NetBSD - a UNIX based operating systems, and OpenWrt  [^34] https://tests.reproducible-builds.org/openwrt/openwrt.html - operating system for embedded devices.
@@ 7.2. Verifying Reproducibility
Outside the Java ecosysem, Malka et al. (Malka et al., 2025) investigate reproducibility in functional package manager Nix. The paper proves that functional package managers help reproduciblity by showing 91% of the packages in Nix are reproducible. Similarly, Bajaj et al. (Bajaj et al., 2023) investigate reproduciblity in Debian and Arch Linux packages. Apart from uncovering the causes of unreproducibility, they also do a survival analysis of packages to see how long it takes to fix the unreproducibility and how long the package stays reproducible. They show that Arch Linux pacakges become reproducible sooner than Debian packages. However, a Debian package is likely to stay reproducible for longer before it becomes unreproducible again. Moritz (Moritz, 2023) proposes a decentralized protocol on Hyperledger Fabric to verify reproducibility. The protocol involves comparing the locally generated buildinfo file with the one on Debian registry. Linderud (Linderud, 2019) also proposes a protocol to verify buildinfo files, but a layer of transparency log is added in order to ensure reproducibility status for specific builds are immutable. Drexel et al. (Drexel et al., 2024) contributes with an independent rebuilder that verifies reproducibility of Arch Linux packages based on this approach. Vu et al. (Vu et al., 2021) focus on finding the most common files and APIs that differ between source repository and Python wheel published on PyPI. Goswami et al. (Goswami et al., 2020) investigate reproducibility in NPM packages. They compare the minified artifacts generated by JavaScript bundlers to report the reasons of unreproducibility. Although there are overlap in taxonomies reported by these works, our analysis focuses on Java ecosystem and proposes novel reasons for unreproducibility. For example, variations in MANIFEST, JVM bytecode, and versioning properties are first seen in our work. Even for reasons such as timestamps, we are the first to report 9 ways that it can be embedded in a Maven release.
@@
Some works analyze single projects rather than the entire ecosystems. Shi et al. (Shi et al., 2022) investigate reproducibility for 4 commercial systems running in Huawei and CentOS - a production grade Linux operating system. Pöll et al. (Pöll and Roland, 2022) discover reproducibility issues while rebuilding Android 5 to 12 images. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) investigate reproducibility for 16 versions of the TrueCrypt encryption tool. Finally, Andersson (Andersson, 2024) investigates reproducibility for Go Ethereum binary. None of the subjects in these related work are Java based. The work of Pöll et al. (Pöll and Roland, 2022) is Java based, but Android images do not compile to JVM bytecode like traditional Java applications. Hence, ours paper is the first one to investigate reproducibility of normal Java applications at scale.
The most closely related work here is the work by Xiong et al. (Xiong et al., 2022) that compares 59 Maven projects on Maven Central and internal registry of Huawei with the artifacts generated by the build process. They discover build reproducibility problems due to environment, JDK, multithreading, and other tools. Our work differs in the scale of the dataset (ours is 4$\times$ larger) and the taxonomy of the root causes of unreproducibility. Our taxonomy has more categories than the one proposed by Xiong et al. incl. new essential categories such as variations in filesystem and software bill of materials.
@@
Benedetti et al. (Benedetti et al., 2025) verify the reproducibility of 12,000 Ruby, PyPI, and Maven packages using this approach. They keep all inputs same except the environment which is changed for each build using reprotest  [^36] https://salsa.debian.org/reproducible-builds/reprotest. reprotest can change locale, timezone, and number of CPUs for example. Ren et al. (Ren et al., 2018, 2022) and Leija et al. (Navarro Leija et al., 2020) propose a technique to localize and fix the root cause of unreproducibility in Java artifacts. They build the source files and dependencies for Debian packages twice and compare the artifacts using diffoscope  [^37] https://diffoscope.org/. Lagnöhed (Lagnöhed, 2024) propose a tool rmake built on top of make with support for building projects twice and comparing the artifacts. These works differ from our work in the approach they take to verify reproducibility. We claim that the first approach uncovers more reasons of unreproducibility as it compares artifacts generated by build process which have maximum differences in environment. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) also emphasizes that there should be maximum difference between the environments to detect unreproducibility.

### v2

@@ 4.3.5. Unreproducible Versioning Properties
Git properties are embedded in Jar files in git.json or git.properties files. Maven plugin git-commit-id-maven-plugin  [^19] https://github.com/git-commit-id/git-commit-id-maven-plugin is commonly used to do so. These properties include timezone of commit, remote URL, number of git tags in repository, name of build host, time of build, name of builder, email of builder, branch name, number of commits, local branch name, and total commits in the repository. All of them may diverge between the builder and rebuilder. For example, value of git.tags vary for Maven release org.apache.drill:drill-opentsdb-storage:1.21.0 due to additional number of tags in the repository during rebuild. Mitigation: The rebuilder should strip out the git properties from the Jar file before comparison or avoid comparison of the Git properties. Another solution is to fix the values for each git property. For example, total commits should be calculated up to the commit that is used to build the artifact. See issue on our repository  [^20] https://github.com/chains-project/reproducible-central/issues/19 to see more examples of differences in the git properties.
@@ 4.3.6. Unreproducible Timestamps
However, we observe that there are 10 other ways timestamps can be embedded in the JAR file. In our dataset, we observe that: 1) they can be embedded in the properties file section˜F.1, 2) generated documentation section˜F.2, 3) shell scripts section˜F.3, 4) executable binaries section˜F.4, 5) software bill of materials section˜F.5, 6) JVM bytecode section˜F.6, 7) file metadata section˜F.7, 8) MANIFEST.MF section˜F.8, and 9) NOTICE section˜F.9. 10) servlets created by Jasper JSP compiler section˜F.10. Mitigation: The solution proposed by Maven is partial, more engineering is needed to support project.build.outputTimestamp in all Maven plugins that create the above mentioned files.
@@ 4.3.7. Miscellaneous Unreproducibility Reasons
We report reasons of reproducibility that are specific to a particular released module. They do not fit into the categories above, so we report them here with suggested mitigations.
Version Ranges Only one release org.apache.aries:org.apache.aries.jax.rs.whiteboard:2.0.2 is observed to use version range for flattened-maven-plugin which affects the generation of the build manifest, pom.xml. It is better to pin exact versions as it contributes to reproducible builds and is also a good security practice (Javan Jafari et al., 2023).
UUIDs: Non-deterministic UUID generation in configuration files prevents reproducibility.
JavaScript Bundling: Bundled JS files can have different variable names and line endings across builds. This has also been reported in the literature (Goswami et al., 2020).
Microbenchmarks: Performance testing results should not be embedded in release artifacts.
Index Files: Cache files from tools like Jandex, EFX Toolkit, and Spotless Formatter are not needed for end users.
@@ 4.3.8. Mysteries
We observe changes in NOTICE, build manifests, and other configuration files that are not due to any of the above reasons and thus, we do not know what caused them and cannot propose a mitigation strategy for them. Hence, we classify them under this section. We discuss the changes one by one in appendix˜H.
4.4. Key Novelty of the Taxonomy
Table 2 emphasizes the fine-grained causes for build unreproducibility that have never been studied in the literature before. Each ✓ indicates a novel finding in our work while empty cells represent causes that have already been reported in the literature, in particular by Xiong et al. (Xiong et al., 2022).
Most of the categories that we have identified are not present in the related work  (Xiong et al., 2022). The high-level reasons for unreproducibility related to SBOM and Filesystem are not mentioned at all in the related work. Other categories are present in the work of Xiong et al., but we identify finer subcategories. For example, we report nine more subcategories for build manifest.
Finally, a fundamental difference between our work and Xiong et al. (Xiong et al., 2022) is that we perform a comparison between reference and rebuild versions of the same artifact while they build twice and compare. We have described these fundamentally different approaches in section˜8.2.
@@ 4.4. Key Novelty of the Taxonomy
We identify 6 main reasons of unreproducibility in Java and their mitigation: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. We also report novelty of our findings in Table 2 with respect to the related work.
@@ 5.2. Methodology
Chains-Rebuild  [^23] https://github.com/chains-project/chains-rebuild/commit/6dd67d5c7ac4db112f3419b5132d8f80a22cbe65 is our own fork of OSS-Rebuild, with crucial improvements dedicated to Maven reproducibility. We add support for canonicalizing build manifests and embedded versioning properties in JAR files as suggested in our taxonomy (Table 2). For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file, a spurious change which is removed by Chains-Rebuild by directly editing the byte buffer of the JAR file. For versioning properties, we remove the git.properties files embedded in the JAR file. Finally, we remove the pom.properties file. These additional canonicalization features have been proposed to the OSS-Rebuild developers through 6 Pull Requests (#339, #360, #363, #377, #414, and #413) .
We run OSS-Rebuild and Chains-Rebuild on the 12,803 reference and rebuild artifact pairs in order to analyze how effectively they make the build reproducible by canonicalization.
@@ 5.3. Results
OSS-Rebuild (4ef4c01) 1,205 (9.41%) 11,598 (90.59%)
Chains-Rebuild (6dd67d5) 3,405 (26.60%) 9,398 (73.40%)
Table 3. Effectiveness of OSS-Rebuild and Chains-Rebuild on mitigating 12,803 unreproducible artifacts via canonicalization.
Table 3 shows the results of OSS-Rebuild and Chains-Rebuild on 12,803 unreproducible artifacts. The first column is the name of the tool and its commit hash we use to canonicalize the artifacts. The second column shows the number of artifact pairs that are successfully canonicalized by OSS-Rebuild and Chains-Rebuild. The third column shows the number of artifact pairs that are not properly canonicalized.
@@
Our publicly available prototype Chains-Rebuild can canonicalize (3,405) 26.60% of the artifacts from our dataset. Our study is the first large-scale experiment on artifact canonicalization, demonstrating that it is a promising mitigation of build unreproducibility..
@@ 6.2. Methodology
Bytecode canonicalization is a process of transforming the bytecode of a program into a representation that is independent of specific implementation details inserted by the compiler. This is either done by converting the bytecode to an intermediate form or by rewriting the bytecode. For example, subtraction of two integers is implemented as addition of a positive and negative integer in Java 6 and higher versions. Meanwhile, in Java 5 and lower versions, it is implemented as subtraction of two integers. Hence, if the program is compiled with Java 5 and 6, the two bytecode representations differ even though the program is semantically the same — performing subtraction of two integers. Canonicalization removes these differences by converting the bytecode to an intermediate form which is free from the implementation details that vary across different versions of the compiler.
@@
We run jNorm version 1.0.0 on the 936 reference rebuild artifact pairs that have differences in the bytecode. For maximum canonicalization, we run the tool with all the command line options as listed in the README 242424jNorm tool — https://github.com/stschott/jnorm-tool/tree/cec4645c5c9b52f73c347349bf14945b0eb55c87 of the repository and tool’s help output. As a result of canonicalizing the artifacts, jNorm produces files in the Jimple (Raja Vallée-Rai and Laurie J. Hendren, 1998) format. We save the canonicalized Jimple files for both the reference and the rebuild artifacts. There are cases where jNorm crashes and canonicalization fails, which we consider as a failure. We do not analyze them further as they are bugs in tool, but we upload the logs for future debugging. We run the Linux diff tool over the canonicalized pairs of Jimple files. If there is no difference, it means that canonicalization is a success, for reproducibility. In other words, successful canonicalization by jNorm produces two identical artifacts. We report the number of cases where the diff between the pair is empty. Finally, we analyze the cases where jNorm is unable to fully canonicalize the bytecode in order to understand its fundamental limitations and identify required future work.
@@ 6.3. Results
Out of 936 artifacts, jNorm’s execution leads to the following distribution: it successfully canonicalizes 275 (29.7%) artifacts which means it produces two identical artifacts; it leaves 508 (53.2%) artifacts with a diff after canonicalization; and there are 153 (17.1%) crashes, which are ignored.
@@
Next, we study the reasons why jNorm fails to produce identical artifacts when run on 508 cases. We categorize them based on the features in the bytecode that are not properly canonicalized by jNorm. Note that some unreproducible artifacts may fall into multiple categories as it is possible jNorm is unable to canonicalize multiple problems in the same artifact.
@@
    Listing 2: jNorm diff for control flow difference.
@@
We note that multiple canonicalization tools can be used in conjunction. For instance, Chains-Rebuild and jNorm can be used together to canonicalize the JVM bytecode and the JAR files of a Maven artifact. Since Chains-Rebuild can canonicalize 3,405 artifacts and jNorm can canonicalize 275 artifacts, we can combine the 28.74% (3,680 / 12,803). Out of 3,680 artifacts, 2,939 artifacts are canonicalized such that 2,493/ 8,292 (30.06%) releases become fully reproducible.
@@
jNorm is able to canonicalize 275 (29.38%) out of 936 bytecode artifacts. However, there are 508 (54.27%) artifacts that are still unreproducible after canonicalization. We present in detail four limitation of bytecode canonicalization related to structural change, control flow, embedded data, and optimization. Finally, our experiments demonstrate that combining artifact and bytecode canonicalization successfully mitigates unreproducibility in 28.74% (3,680 / 12,803) artifacts.
7. Best Practices for Builders and Rebuilders
7.1. Builders
Builders can improve reproducibility by ensuring deterministic build inputs, such as pinning dependency versions (), clearing caches before deployment, and using standard build commands rather than ad hoc scripts. They should avoid embedding non-deterministic data like branch names, timestamps, usernames, or UUIDs in manifests and properties, and instead use plugins that enforce deterministic ordering. To support verification, builders should also document and publish the exact JDK version and vendor used for releases, and generate SBOMs in a consistent and reproducible manner.
7.2. Rebuilders
Rebuilders should strive to mirror builder environments as closely as possible by using the same JDK vendor and version, along with consistent settings for locale, time zone, and operating system. When unreproducibility persists, they can apply canonicalization tools such as OSS-Rebuild and jNorm to remove spurious differences in bytecode and artifacts. These tools are carefully crafted to be able to canonicalize unreproducibility issues in the bytecode and artifacts. The remanining diff after first canonicalization can help as a feedback to the builder to fix the unreproducibility issue or to write more canonicalization rules.
@@ 8.1. Reproducibility in Applications
The Reproducible Builds  [^30] https://reproducible-builds.org/ project works on the reproducibility of coreboot  [^31] https://tests.reproducible-builds.org/coreboot/coreboot.html — an open-source firmware for x86 and ARM systems, Debian  [^32] https://tests.reproducible-builds.org/debian/reproducible.html — a Linux distribution, FreeBSD  [^33] https://tests.reproducible-builds.org/freebsd/freebsd.html and NetBSD — a UNIX based operating systems, and OpenWrt  [^34] https://tests.reproducible-builds.org/openwrt/openwrt.html — operating system for embedded devices.
@@ 8.2. Verifying Reproducibility
Outside the Java ecosysem, Malka et al. (Malka et al., 2025) investigate reproducibility in functional package manager Nix. The paper proves that functional package managers help reproduciblity by showing 91% of the packages in Nix are reproducible. Similarly, Bajaj et al. (Bajaj et al., 2023) investigate reproduciblity in Debian and Arch Linux packages. Apart from uncovering the causes of unreproducibility, they also do a survival analysis of packages to see how long it takes to fix the unreproducibility and how long the package stays reproducible. They show that Arch Linux packages become reproducible sooner than Debian packages. However, a Debian package is likely to stay reproducible for longer before it becomes unreproducible again. Moritz (Moritz, 2023) proposes a decentralized protocol on Hyperledger Fabric to verify reproducibility. The protocol involves comparing the locally generated buildinfo file with the one on Debian registry. Linderud (Linderud, 2019) also proposes a protocol to verify buildinfo files, but a layer of transparency log is added in order to ensure reproducibility status for specific builds are immutable. Drexel et al. (Drexel et al., 2024) contributes with an independent rebuilder that verifies reproducibility of Arch Linux packages based on this approach. Vu et al. (Vu et al., 2021) focus on finding the most common files and APIs that differ between source repository and Python wheel published on PyPI. Goswami et al. (Goswami et al., 2020) investigate reproducibility in NPM packages. They compare the minified artifacts generated by JavaScript bundlers to report the reasons of unreproducibility. Although there are overlap in taxonomies reported by these works, our analysis focuses on Java ecosystem and proposes novel reasons for unreproducibility. For example, variations in MANIFEST, JVM bytecode, and versioning properties are first seen in our work. Even for reasons such as timestamps, we are the first to report 9 ways that it can be embedded in a Maven release.
@@
Some works analyze single projects rather than the entire ecosystems. Shi et al. (Shi et al., 2022) investigate reproducibility for four commercial systems running in Huawei and CentOS — a production grade Linux operating system. Pöll et al. (Pöll and Roland, 2022) discover reproducibility issues while rebuilding Android 5 to 12 images. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) investigate reproducibility for 16 versions of the TrueCrypt encryption tool. Finally, Andersson (Andersson, 2024) investigates reproducibility for Go Ethereum binary. None of the subjects in these related work are Java based. The work of Pöll et al. (Pöll and Roland, 2022) is Java based, but Android images do not compile to JVM bytecode like traditional Java applications. Hence, ours paper is the first one to investigate reproducibility of normal Java applications at scale.
@@
Benedetti et al. (Benedetti et al., 2025) verify the reproducibility of 12,000 Ruby, PyPI, and Maven packages using this approach. They keep all inputs same except the environment which is changed for each build using reprotest  [^36] https://salsa.debian.org/reproducible-builds/reprotest. reprotest can change locale, timezone, and number of CPUs for example. Ren et al. (Ren et al., 2018, 2022) and Leija et al. (Navarro Leija et al., 2020) propose a technique to localize and fix the root cause of unreproducibility in Java artifacts. They build the source files and dependencies for Debian packages twice and compare the artifacts using diffoscope  [^37] https://diffoscope.org/. Lagnöhed (Lagnöhed, 2024) propose a tool rmake built on top of make with support for building projects twice and comparing the artifacts.
In Java, the most closely related work here is the work by Xiong et al. (Xiong et al., 2022). They compare 59 Maven projects on Maven Central and internal registry of Huawei with the artifacts generated by the build process. They discover build reproducibility problems due to environment, JDK, multithreading, and other tools. Our work differs in the scale of the dataset (ours is 4$\times$ larger) and the taxonomy of the root causes of unreproducibility. Our taxonomy has more categories than the one proposed by Xiong et al. including new essential categories such as variations in filesystem and software bill of materials.
All of these works differ from our work in the approach they take to verify reproducibility. We claim that the first approach uncovers more reasons of unreproducibility as it compares artifacts generated by build process which have maximum differences in environment. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) also emphasizes that there should be maximum difference between the environments to detect unreproducibility.